For December 31, 20X1, the balance sheet of Baxter Corporation was as follows:
  

Current Assets

Liabilities

Cash

$

19,000

Accounts payable

$

21,000

Accounts receivable

24,000

Notes payable

29,000

Inventory

34,000

Bonds payable

59,000

Prepaid expenses

12,900

Fixed Assets

Stockholders’ Equity
Gross plant and equipment

$

259,000

Preferred stock

$

29,000

Less: Accumulated depreciation 51,800 Common stock

64,000

Paid in Capital

34,000

Net plant and equipment

$

207,200

Retained earnings

61,100

Total assets

$

297,100

Total liabilities and stockholders’ equity

$

297,100


Sales for 20X2 were $265,000, and the cost of goods sold was 55 percent of sales. Selling and administrative expense was $26,500. Depreciation expense was 12 percent of plant and equipment (gross) at the beginning of the year. Interest expense for the notes payable was 8 percent, while the interest rate on the bonds payable was 16 percent. This interest expense is based on December 31, 20X1 balances. The tax rate averaged 40 percent.

$2,900 in preferred stock dividends were paid, and $6,086 in dividends were paid to common stockholders. There were 10,000 shares of common stock outstanding.

During 20X2, the cash balance and prepaid expenses balances were unchanged. Accounts receivable and inventory increased by 8 percent. A new machine was purchased on December 31, 20X2, at a cost of $44,000.

Accounts payable increased by 20 percent. Notes payable increased by $6,900 and bonds payable decreased by $14,500, both at the end of the year. The preferred stock, common stock, and capital paid in excess of par accounts did not change.

a. Prepare an income statement for 20X2. (Round EPS answer to 2 decimal places.)

Baxter Corporation
20X2 Income Statement
Earnings available to common stockholders
Shares outstanding
Earnings per share


  b. Prepare a statement of retained earnings for 20X2.

Baxter Corporation
20X2 Statement of Retained Earnings
Retained earnings balance, January 1, 20X2
Retained earnings balance, December 31, 20X2

c. Prepare a balance sheet as of December 31, 20X2. (Amounts to be deducted should be indicated with parentheses or a minus sign.)

Baxter Corporation
20X2 Balance Sheet
Assets Liabilities and Stockholders’ Equity
Current Assets:
Total current assets Total liabilities
Stockholders' Equity
Fixed assets:
Net plant and equipment
Total stockholders' equity
Total assets Total liabilities and stockholders' equity
Answer #1

Answer a:

Baxter Corporation 20X2 Income Statement Calculations: $265,000 $145,750 $119,250 $31,080 $26,500 $61,670 $11,760 $49,910 $19,964 $29,946 $2,900 $27,046 10,000 $2.70 $6,086 Sales Cost of Goods sold :55% * 265000 Gross Income Deprecation -12% *259000 Selling and administrative expense EBIT Interest expense % * 29000 + 16% *59000 EBT Tax at 40% Net Income Less, Preference stock dividends Earnings available to common stockholders Shares outstandin Earnings per share Dividends paid to common stockholders =40% * 499 10

Answer b:

Baxter Corporation 20X2 Statement of Retained Earnings Retained earnings balance, January 1, 20X2 Net Income livicleicl paicl $61,100 $29,946 $8,986

Dividends paid = Preference stock dividends + Dividends paid to common stockholders = 2900 + 6086 = $8,986

Answer c:

Baxter Corporation 20X2 Balance Sheet Liabilities and Stockholders Equit Assets Current Assets: Cash Account Receivable Invento Prepaid expenses Total current assets 19,000 Accounts payable 25,920 Notes payable 36,720 Bonds payable 12,900 94,540 Total liabilities $25,200 35,900 $ 44,500 S105,600 Stockholders Equity Fixed assets: Less, Accumulated depreciation 303,000 Preferred stock 82,880 Common stock Paid in capital $29,000 $64,000 $34,000 $ 82,060 $209,060 314,660 Total liabilities and stockholdersequity 314,660 Net plant and equipment 220,120 Retained earnings Total stockholders equity Total assetsWorkings:

Accounts receivable = 24000 * (1+8%) = $25,920

Inventory = 34000* (1+8%) = $36,720

Fixed Assets =259000+44000 = $303,000

Accumulated depreciation = 51800 + 31080 =$82,880

Accounts payable =21000*(1+20%) = $25,200

Notes payable =29000 +6900 = $35,900

Bonds payable = 59000-14500 =$44,500